OpenStudy (anonymous):

2^-3 need help with this problem and also 10x^8/5x^-14 (integer exponents and scientific notations)

OpenStudy (anonymous):

Just remember that when you multiply exponents with the same base you add their exponents. When you divide you subtract \[a^b \times a^c = a^{b+c}\] \[\frac{a^b}{a^c} = a^{b-c}\]

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

what do u mean by the same base?

OpenStudy (anonymous):

Well, in the above example I gave, the base is a. It's the thing you are raising to a power.

OpenStudy (anonymous):

answer 2x^6